OpenStudy (anonymous):

solve in R the equations : cos(2x)=√2/2 cos(3x)=sin(2x) tan^2(x)-((√3) +1)tan(x)+√3=0

OpenStudy (anonymous):

1. cos 2x = V2/2 = cos Pi/4 , and = cos 7Pi/4 a)2x = Pi/4 --> x = Pi/2 b) 2x = Pi/4 -> x = 7Pi/8 2. cos 3x = sin 2x = cos (Pi/2 - 2x) , and = cos (2Pi - Pi/2 + 2x) a) 3x = Pi/2 - 2x -> 5x = Pi/2 --> x = Pi/10 b) 3x = 3Pi/2 + 2x -> x = 3Pi/2

OpenStudy (anonymous):